can you please explain this to me?
Sole the system of equations using eh addition (elimination) method. If the answer is a unique solution, present it as an ordered pair: (x.y). If not, specify whether the answer is “no solution” or “ Inifinitely many solutions.”
9x + 2y = 2
3x + 5y = 5
thank you
julie

Answer by jgr45(22)

9x + 2y = 2
3x + 5y = 5
Start by multiplying the bottom by -3 (the top stays as is):
9x + 2y = 2
-9x - 15y = -15
Combine (x's cancel):
-13y = -13
y = 1
Plug 1 for y in either equation and solve for x.
3x + 5(1) = 5
3x + 5 = 5
3x = 0
x = 0
Point of intersection = (0,1)